[Ncep.nhc.nco_contacts] TSB Morning Rounds - Fri Sep 15, 2017

Christopher Mello - NOAA Federal christopher.mello at noaa.gov
Fri Sep 15 17:20:11 UTC 2017


All,
There  is a wxconnect problem that stopped the GOESR data from flowing. .
AWC, NHC, SPC all stopped receiving product via LDM at 1430Z.

The SDM is  working to resolve the issue...

Chris



On Fri, Sep 15, 2017 at 12:27 PM, TSB Admin <nhc.tsbadmin at noaa.gov> wrote:

> Here is today's summary of NHC computer operations:
> --- Matt Onderlinde
>
> ----------
>
> *TAFB:*
>
> 1) Security camera display remains out in WFO Miami and TAFB ops areas
> (see previous log entries for request for training on how to fix this for
> TAFB/WFO leads)
>
> 2) GOES-E 2345z image did not come in (data drop out?). This impacted
> classifications for Jose and Max.
> *HSU:*
>
> none reported
> *Other items:*
>
> n/a
> *NHC RFCs:*
>
> n/a
>
>


-- 
Chris Mello
National Hurricane Center
TSB
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170915/888c4cc1/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list